# Tunable Topological Phases in Quantum Kirigamis

### v1

@@ I Wave packet evolution and estimation of $r(\theta),t(\theta)$
We study the junction by first choosing two finite sized square lattices (describing two Kirigami blocks) each of size $L\times L$ such that the total number of sites are $N=2L^{2}$. Every site has two orbitals $A,B$ such that the total number of fermionic orbitals is $2N$. The Hamiltonian describing these two blocks separately is $H_{\text{BHZ}}$ and that describing the couplings between these two blocks is $\equiv H_{\text{coupling}}$. Thus the total Hamiltonian of the system is given by,
$H=H_{\rm{BHZ}}+H_{\rm{coupling}}$ (S1)
@@
$H_{\text{BHZ}}=\sum_{i,\delta}\Psi^{\dagger}_{i}T_{\delta}\Psi_{i+\delta}+\text{h.c.}+\sum_{i}\Psi^{\dagger}_{i}\Gamma\Psi_{i}$ (S2)
where $\delta=\hat{x},\hat{y}$, the unit vectors in the $x$ and $y$ directions defined with respect to a local axis defined for every Kirigami block. $\Psi_{i}\equiv\Big{(}c_{iA},c_{iB}\Big{)}^{T}$ where $c_{iA},c_{iB}$ are the fermionic annihilation operators corresponding to the two orbitals $A,B$ at site $i$. $\Gamma$ describes the onsite energies and $T_{\hat{x}}$ and $T_{\hat{y}}$ are the hopping matrices given by:
$T_{\hat{x}}=-\frac{1}{2}(\sigma_{z}+i\sigma_{x}),T_{\hat{y}}=-\frac{1}{2}(\sigma_{z}+i\sigma_{y}),\Gamma=(M+2)\sigma_{z}$ (S3)
In a translationally invariant system, this Hamiltonian in the momentum space gives rise to eqn. (1) in the main text. Note that as the angle $\theta$ is changed, such that the local coordinate system of every Kirigami block obtains a relative angle with respect to another Kirigami block, the hopping amplitudes within any Kirigami block does not change. For instance as shown in Fig. S1 the relative angle between the two Kirigami blocks with local coordinates $(x,y)$ and $(x^{\prime},y^{\prime})$ respectively has a mutual angle of $2\theta$.
Figure S1: Model: (a) A junction is modelled with two square Kirigami blocks meeting at a node. The interblock couplings are limited to a width $w$ and has a angle dependent hopping of strength $\lambda(\theta)$. (b) While each block has a size $L\times L$, it has a local coordinate axis which rotates as a function of deformation angle. The Kirigami blocks can be divided into four regions I-IV to estimate the scattering matrix coefficients (see text).
$H_{\text{coupling}}$: Next we couple the two Kirigami blocks with interblock hoppings as follows. Labelling the sides (edges) forming the junction ($s_{e}$=1-4) as shown in Fig. S1(a), we identify the sites along the edges $n=1,2,\ldots w$. Thus $w$ number of sites, denoting the width of the edge, couple between the sides facing each other. For instance in Fig. S1(a), $s_{e}=1$ and $s_{e}=4$, and similarly $s_{e}=2$ and $s_{e}=3$. Labeling the fermionic creation operators on the edge $s_{e}$ at site $n$ as $\Psi^{\dagger}_{n,s_{e}}$, the $H_{\rm{coupling}}$ is therefore given by
@@
$\lambda(\theta)=\alpha\exp(-\beta d(\theta))$ (S5)
where $\alpha$ and $\beta$ are tuning parameters. We set $\alpha=1,\beta=1$ unless otherwise stated. The choice of parameters and the coupling Hamiltonian have been kept such that at $\theta=0$, $\lambda(\theta)=1$ thus the coupling reduces to the bare hopping scale of the rest of the BHZ system.
Having described the Hamiltonian of the system, we next discuss the wavepacket evolution and estimation of $r(\theta)$ and $t(\theta)$. Given a choice of $M$ and $\theta$, we first diagonalize $H$ (see eqn. (S1)) to obtain the complete eigenspectrum. In the gapped topological regime of the system, the bulk gap is given $\Delta_{g}(M)$. We form a projector out of the single particle wavefunctions $|m\rangle$ such that their eigen energies $\epsilon_{m}$ are $|\epsilon_{m}|<\Delta_{g}(M)$. The projector is thus given by
$\hat{P}=\sum_{m,|\epsilon_{m}|<\Delta_{g}(M)}|m\rangle\langle m|$ (S6)
@@
Figure S2: System size ($N$) dependence: (a) $r(\theta)$ and $t(\theta)$ in BHZ junction modelling (see Fig.2(a) of main text) when $N=450$ and $N=1250$ with $\alpha=1,\beta=1,w=5$ and $M=-3.6$. (b) $r(\theta)$ and $t(\theta)$ in Non-Hermitian junction modelling (see Fig.2(c) of main text) when $N=400$ and $N=560$ with $\alpha=1.5,\beta=0.25$ and $w=5$.
Figure S3: Coupling dependence:  $r(\theta)$ and $t(\theta)$ in BHZ junction modelling (see Fig.2(a) of main text) with ($N=1250$, $M=-3.6,\beta=1$ and $w=5$). (a) $\alpha=0.2$ corresponds to the $r\approx 1$ and $t\approx 0$. (b) For $\alpha=0.4$, $r=t$ when $\theta\approx 0$ or $\frac{\pi}{2}$. (c) For $\alpha=1.2$, $r=t$ when $\theta\approx\frac{\pi}{4}$. (d) For $\alpha=1.5$ $t(\theta)>r(\theta)$.
Figure S4: $w$ dependence: The junction width $(w)$ dependence of $r(\theta)$ and $t(\theta)$ in BHZ junction modelling (see Fig.2(a) of main text) with ($N=1250$, $\alpha=1,\beta=1$ and $M=-3.6$)
To start the wavepacket evolution we initialize a unit state on the top corner of the side 1 on the A orbital. We project this state using the projector above (see eqn. (S6)) and renormalize it to obtain the initial state $|\psi_{\text{in}}\rangle$. This state is further evolved using the Hamiltonian (see eqn. (S1)) to a later time such that final state $|\psi_{\text{fi}}\rangle$ at time $t$ is given by
$|\psi_{\text{fi}}\rangle=\exp(-iHt)|\psi_{\text{in}}\rangle$ (S7)
The initial packet moves towards the junction and scatters into edges 2 and 4. Given the unidirectional chiral edge states, there is neither any backscattering nor any amplitude on edge 3. After a time when the packet has scattered from the junction, we calculate the total probability density on the sides 2 and 4. To estimate the reflected probability density we evaluate
$p_{2}=\sum_{i\in\rm{region~{}II},\gamma={A,B}}|\langle i\gamma|\psi_{\text{fi}}\rangle|^{2}$ (S8)
@@
$p_{4}=\sum_{i\in\rm{region~{}IV},\gamma={A,B}}|\langle i\gamma|\psi_{\text{fi}}\rangle|^{2}$ (S9)
The various regions (I-IV) are shown in Fig. S1. Consistent with the expectations we find that the corresponding quantities in regions III and I $p_{3}=p_{1}=0$ after the scattering event. Having estimate $p_{2}$ and $p_{4}$, we calculate $r(\theta)=\sqrt{p_{2}}$ and $t(\theta)=\sqrt{p_{4}}$. This procedure is repeated for various values of $M$ and $\theta$ to obtain the variations of $r(\theta),t(\theta)$ and the numerically obtain the phase diagrams as reported in the main text. The procedure remains the same for the triangular motif as well, where the Kirigami block is still a square lattice BHZ but the boundaries are formed into a triangular motif. Other parameters and the coupling Hamiltonian are kept the same.

### v2

@@ V.1 Wave-packet dynamics
We study the junction by first choosing two finite-sized square lattices (describing two Kirigami blocks) each of size $L\times L$ such that the total number of sites is $N=2L^{2}$. Every site has two orbitals $A$, $B$ such that the total number of fermionic orbitals is $2N$. The Hamiltonian describing these two blocks separately is $H_{\text{BHZ}}$ and that describing the couplings between these two blocks is $\equiv H_{\text{coupling}}$. Thus, the total Hamiltonian of the system is given by,
$H=H_{\rm{BHZ}}+H_{\rm{coupling}}$ (4)
Figure 3: Model: (a) A junction is modelled with two square Kirigami blocks meeting at a node. While each block has a size $L\times L$, it has a local coordinate axis that rotates as a function of the deformation angle. The Kirigami blocks can be divided into four regions I-IV to estimate the scattering matrix coefficients (see section V.1). (b) Schematic showing that the interblock couplings are limited to a width $w$ and have an angle-dependent hopping of strength $\lambda(\theta)$ between sites separated by a distance $d(\theta)$. (c) Schematic showing that an incident electronic wave packet (from edge channel 1) on a node, gets reflected (on channel 2) and transmitted (on channel 4) by scattering matrix elements $r$ and $t$ respectively. (d) Behavior of $r(\theta),t(\theta)$ as a function of $\theta$ using wave-packet dynamics (see section V.1) for $M=-3.6,\alpha=1,\beta=1,w=5,N=1250$. Here $N$ is the total number of sites in both the blocks forming the junction. Further, using these obtained $r(\theta),t(\theta)$ for the ENJH model, where the node itself is replaced by four sites, with inter-site non-Hermitian hoppings $r$ and $t$. Further, using the wavepacket dynamics for the ENHJ model, the plot $r(\theta),t(\theta)$ exactly match with those obtained using the BHZ model (see section V.2).
@@
$H_{\text{BHZ}}=\sum_{i,\delta}\Psi^{\dagger}_{i}T_{\delta}\Psi_{i+\delta}+\text{h.c.}+\sum_{i}\Psi^{\dagger}_{i}\Gamma\Psi_{i}$ (5)
where $\delta=\hat{x},\hat{y}$, the unit vectors in the $x$ and $y$ directions defined with respect to a local axis defined for every Kirigami block (see Fig. 3(a)). $\Psi_{i}\equiv\Big{(}c_{iA},c_{iB}\Big{)}^{T}$ where $c_{iA},c_{iB}$ are the fermionic annihilation operators corresponding to the two orbitals $A,B$ at site $i$. $\Gamma$ describes the onsite energies and $T_{\hat{x}}$ and $T_{\hat{y}}$ are the hopping matrices given by:
$T_{\hat{x}}=-\frac{1}{2}(\sigma_{z}+i\sigma_{x}),T_{\hat{y}}=-\frac{1}{2}(\sigma_{z}+i\sigma_{y}),\Gamma=(M+2)\sigma_{z}$ (6)
where $\sigma_{x},\sigma_{y},\sigma_{z}$ are the Pauli matrices. In a translationally invariant system, this Hamiltonian in the momentum space gives rise to eqn. (1). Note that as the angle $\theta$ is changed, such that the local coordinate system of every Kirigami block obtains a relative angle with respect to another Kirigami block, the hopping amplitudes within any Kirigami block does not change. For instance as shown in Fig. 3(a) the relative angle between the two Kirigami blocks with local coordinates $(x,y)$ and $(x^{\prime},y^{\prime})$ respectively has a mutual angle of $2\theta$.
$H_{\text{coupling}}$: Next we couple the two Kirigami blocks with interblock hoppings as follows. Labeling the sides (edges) forming the junction ($s_{e}$=1-4) as shown in Fig. 3(b), we identify the sites along the edges $n=1,2,\ldots w$. Thus $w$ number of sites, denoting the width of the junction, are coupled between the sides facing each other (for instance, in Fig. 3(b), $s_{e}=1$ and $s_{e}=4$ and similarly $s_{e}=2$ and $s_{e}=3$). Labeling the fermionic creation operators on the edge $s_{e}$ at site $n$ as $\Psi^{\dagger}_{n,s_{e}}$, the $H_{\rm{coupling}}$ is therefore given by
@@
$\lambda(\theta)=\alpha\exp(-\beta d(\theta))$ (8)
where $\alpha$ and $\beta$ are tuning parameters. We set $\alpha=1,\beta=1$ unless otherwise stated. The choice of parameters and the coupling Hamiltonian has been kept such that at $\theta=0$, $\lambda(\theta)=1$ thus the coupling reduces to the bare hopping scale of the rest of the BHZ system.
Having described the Hamiltonian of the system, we next discuss the wavepacket evolution and estimation of $r(\theta)$ and $t(\theta)$. Given a choice of $M$ and $\theta$, we first diagonalize $H$ (see eqn. (4)) to obtain the complete eigenspectrum. In the gapped topological regime, the bulk gap in the periodic boundary condition is given $\Delta_{g}(M)$. We form a projector out of the single particle wavefunctions $|m\rangle$ such that their eigen energies $\epsilon_{m}$ are $|\epsilon_{m}|<\Delta_{g}(M)$. The projector is thus given by
$\hat{P}=\sum_{m,|\epsilon_{m}|<\Delta_{g}(M)}|m\rangle\langle m|$ (9)
@@
Figure 4: Topological mechanical insulator:  The time evolution of a wavepacket under the ENHJ formalism (see section. V.2) for three configurations of the Kirigami networks at (a) $\theta=15^{\circ}$ for three different times $\tau=\{2,60,120\}$ (shown by arrows). (b) $\theta=40^{\circ}$ and $\tau=\{2,70,90\}$ (c) $\theta=85^{\circ}$ and $\tau=\{2,90,180\}$. While (a) and (c) show that the network is topological (edge states), in (b) one finds that the network is trivial.
To start the wavepacket evolution we initialize a unit state on the top corner of side 1 on the A orbital. We project this state using the projector above (see eqn. (9)) and renormalize it to obtain the initial state $|\psi_{\text{in}}\rangle$. This state is further evolved using the Hamiltonian (see eqn. (4)) to a later time such that final state $|\psi_{\text{fi}}\rangle$ at time $\tau$ is given by
$|\psi_{\text{fi}}\rangle=\exp(-iH\tau)|\psi_{\text{in}}\rangle$ (10)
The initial packet moves towards the junction and scatters into edges 2 and 4. Given the unidirectional chiral edge states, there is neither any back-scattering nor any amplitude on edge 3. After a time when the packet has scattered from the junction, we calculate the total probability density on sides 2 and 4. To estimate the reflected probability density we evaluate
$p_{2}=\sum_{i\in\rm{region~{}II},\gamma={A,B}}|\langle i\gamma|\psi_{\text{fi}}\rangle|^{2}$ (11)
@@
$p_{4}=\sum_{i\in\rm{region~{}IV},\gamma={A,B}}|\langle i\gamma|\psi_{\text{fi}}\rangle|^{2}$ (12)
The various regions (I-IV) are shown in Fig. 3(a). Consistent with our expectations we find that the corresponding quantities in regions III and I are zero ($p_{3}=p_{1}=0$) after the scattering event. Having estimate $p_{2}$ and $p_{4}$, we calculate $r(\theta)=\sqrt{p_{2}}$ and $t(\theta)=\sqrt{p_{4}}$.
Choosing $M=-3.6$ in the topological phase with Chern number $C=-1$ we evaluate $r,t$ for all values of $\theta$. We find an intriguing dependence, where $t(\theta)\sim 1$ near both $\theta\sim 0$ and $\sim\frac{\pi}{2}$ while $r(\theta)>t(\theta)$ near $\theta\sim\frac{\pi}{4}$ (Fig. 3(d)). We have checked that this behavior is independent of the system size (see Appendix A). The choice of parameter values, in particular $w\sim 5$ is kept to ensure a competition with the localization length $\xi$ which changes as a function of $M$. A large value of $w\gg\xi$ will ensure $t\sim 1$ at all values of $\theta$. We also note that the exact values depend on the choice of $\lambda(\theta),w$, and $M$, however, the qualitative behavior of these coefficients do not (see Appendix A). With this the first step of coarse-graining (see Fig. 2) is complete and we move to the ENHJ formalism to investigate the properties of the full network.
V.2 Effective non-Hermitian junction
In order to study the transport property of the full network - we devise an ENHJ formalism where the scattering properties of every node are replaced by a non-Hermitian scatterer as seen in Fig. 2(b). This effectively ensures no back-scattering and with a given value of $r,t$ exactly reproduces the reflection and transmission probabilities through the junction (see Fig. 3(d)). For the complete network, a non-Hermitian Hamiltonian is thus given by
$H=H_{\text{edge}}+H_{\text{hyb}}$ (13)
where
$H_{\text{edge}}=\sum_{i,s_{e}}-c^{\dagger}_{\bm{r}_{i}+\hat{e},s_{e}}c_{\bm{r}_{i},s_{e}}$ (14)
specifies the edge Hamiltonians. $s_{e}$ specifies any edge with the chiral current in direction $\hat{e}$ and $\bm{r}_{i}$ are the discrete positions labeling the sites on any edge. $H_{\text{hyb}}$ represents the mixing between edge modes at two types of nodes $n\in\{N_{1},N_{2}\}$ (having opposite arrow directions, as shown in Fig. 2(b), with yellow and blue colors respectively), such that
$H_{\text{hyb}}$ $=\sum_{n\in N_{1}}\Big{(}[-rd^{\dagger}_{2,n}+td^{\dagger}_{4,n}]d_{1,n}+[td^{\dagger}_{2,n}+rd^{\dagger}_{4,n}]d_{3,n}\Big{)}$ (15)
$+$ $\!\!\sum_{n\in N_{2}}\Big{(}[rd^{\dagger}_{3,n}-td^{\dagger}_{1,n}]d_{2,n}+[td^{\dagger}_{3,n}+rd^{\dagger}_{1,n}]d_{4,n}\Big{)}$
where $s_{e}=$1-4 are the edges labeled anticlockwise at any node, and edges $\{1,3\}$ and $\{2,4\}$ have the incoming (outgoing) and outgoing (incoming) edge currents in nodes $N_{1}(N_{2})$. $d^{\dagger}_{s_{e},n}$ represents the fermionic operators (identical to $c^{\dagger}$ operator) of the edge $s_{e}$ which forms part of the four-site junction, with $r,t$ being the scattering amplitudes of any node $n$. Given the complete morphology of the Kirigami network, every node, in general, can have different values of $r,t$ decided by local junction parameters such as $\theta,w$ etc.
With this coarse-graining step, we can now pose how the complete network behaves given a value of $r,t$ for each node. To make progress we first study the case where $r_{n},t_{n}=(r,t)$ independent of node position, thus leading to a translational symmetry on the network. We first study the Kirigami in the relatively little strain ($\theta=15^{\circ}$) which corresponds to $r,t=(0.52,0.85)$. We again release a wavepacket on the edge of the network and evolve using the ENHJ Hamiltonian (see eqn.13), we find that the wavepacket evolves and moves through the boundary of the network reflecting the topological character of the full network consistent with topologically non-trivial Kirigami blocks (see Fig. 4(a)). When the Kirigami is now stretched to an angle $\theta\sim 40^{\circ}$ with $r,t=(0.78,0.63)$, we notice something surprising. A similar wave packet now scatters into the bulk showing no protection to the edge state, thus suggesting that the complete network is trivial even though each of the blocks is topological (see Fig. 4(b)). A further increase in $\theta=85^{\circ}$ again makes the network topological (see Fig. 4(c)). This reflects the central result of this work, where a mechanical deformation leads to a drastic change in the transport properties of the system. We have further checked that realistic gradients in strain even up to $\sim 15-30\%$ or so do not change these properties (see Appendix. B). The above analysis also suggests the existence of a critical angle, at which the network must transit from a topological to a trivial phase. To uncover such a transition we do further coarse-graining and analyze the system within a Chalker-Coddington formalism.
V.3 Chalker Coddington networks
Having defined the property of every junction and the observation of strain dependent transport of the full network we now pose if there is a critical angle where the transition from topological to trivial phase occurs for the complete Kirigami network. A Chalker-Coddington framework is particularly useful to analyze this[27]. Briefly, the network can be studied within a transfer matrix formalism where an initial state evolves unitarily as [33]
$|\Psi_{f}\rangle=U_{\text{eff}}|\Psi_{i}\rangle$ (16)
where the effective evolution operator, $U_{\text{eff}}$ is decided by the scattering matrix structure of the network. The $U_{\text{eff}}$ is constructed using a median lattice where each edge of the Kirigami block is replaced by a site such that the system has a 16-site unit cell. The ‘hopping’ elements are represented by $r(\theta)$ and $t(\theta)$ with appropriate convention depending on the direction of the edge current. For analytic control we assume that the network is translationally invariant and see if it retains any edge states we set it up on a ribbon geometry which is finite (periodic) in the $y$ ($x$) direction. We find the eigenvalues of $U_{\text{eff}}$ which are of the form $e^{-i\epsilon}$ and plot $\epsilon$ as a function of $k_{x}$. Note that $r,t$ enter as parameters in the transfer matrix. The network description closely follows the formalism developed in Charles et. al.[33], done in the context of Chalker-Coddington networks for integer quantum Hall transition.
Figure 6: Edge state profile:  $P_{y}$ as function of $y$ when $L_{y}=16,r=0.4,t=0.916$. For (I) $k_{x}=-0.5,\epsilon=0.125$ and for (II) $k_{x}=0.5,\epsilon=0.125$
When $r<t$ and $\theta<\theta_{1c}$, we find linearly dispersing edge states on the boundary (see Fig. 5(d)) reflecting that the topological character of the individual Kirigami block also leads to an overall chiral edge in the system (Fig. 5(a)). The electronic probability density $P_{y}$ (as a function of $y$) for given values of $k_{x}$ and $\epsilon$ is shown in Fig. 6. As can be seen, the right (left) moving edge state resides on the bottom (top) of the ribbon. These states are absent when periodic boundary conditions are used in the $y$ direction. However now as $\theta$ is increased, or the structure opens up, the individual scattering elements ($r(\theta),t(\theta)$) modulate approaching a critical point $\theta_{1c}$ where $r=t$. At this point, the system undergoes a phase transition where the $\epsilon$ shows a Dirac-like gap closing. After $\theta>\theta_{1c}$ the gap opens up again (Fig. 5(e)) with no chiral mode on the complete network. This reflects that the Kirigami network has become a “trivial” insulator (see Fig. 5(b)). This is an example of a mechanical deformation-induced topological phase transition. In order to corroborate the results from ENHJ formalism and CC networks we compare the spectrum and show in Appendix. C. With further increase in $\theta$ another critical point $\theta_{2c}$ is approached where the system transits back from the trivial to topological phase (see Fig. 5(c) and (f)). The specific values of $\theta_{1c}$ and $\theta_{2c}$ are dependent on both the microscopic parameters of the Hamiltonian governing each block, as well as the nature of the network. For instance in the case of the square network shown in Fig. 1(a), $\theta_{1c}$ and $\theta_{2c}$ are symmetric about $\pi/4$. For every value of $M$ numerically evaluated $r=t$ points leads to the phase diagram as shown in Fig. 1(c),(d). While we have chosen $r,t$ to be real, a uniform phase does not change $\theta_{1c},\theta_{2c}$. However relative scattering phases either uniform or randomly varying from block to block can lead to both changes in critical angles and new phases [33] which we do not explore here.
We note that this analysis and formalism is only applicable when each of the Kirigami blocks resides in the topological regime i.e. $-4<M<0$. For other values of $M$ when each of the blocks is itself an insulator - mechanical deformation doesn’t render it topological. Each block has a topological phase transition within $-4<M<0$ at $M=0,-2$, and $-4$ where the system contains Dirac cones. Here again, the blocks are themselves metallic and this analysis breaks down. Given the $M=-2$ point sits between the two topological Chern insulator phases of Chern number ±$\pm 1$, $r=t$ by symmetry reasons. Overall the phase diagram is further symmetric about $M=-2$ since apart from the direction of the chiral edge states the physics remains the same.
It is interesting to point out the physics of the $r=t$ line in this system, particularly near $\theta\sim 0$. Given when each Kirigami block is in the topological regime - there exists an edge state with an edge localization length $\xi$, which is dependent on the bulk gap which in turn depends on $M$. Thus whether an edge wavepacket can transmit or reflect from the junction depends on whether $\xi\ll w$ or $\xi\gg w$ respectively. Therefore given a choice of $w$, there exists a critical value of $M$ even within the topological phase, defined as $M^{*}$ at which $\xi(M)\sim w$, therefore leading to a point where $r=t$. This value of $M^{*}$ is thus independent of the system size of the Kirigami block but only dependent on the value of $\theta$ and the parameters defining the junction (see Appendix D). We further note tuning $\alpha,\beta$ which defines the strength of the junction can further tune the values of $r,t$ (see Appendix A).
VI Triangular Network
In order to test these physics in another network, we set up the same problem however now on a Kirigami network where the network deforms between the triangular lattice to a Kagome-like network as a function of deformation angle $\theta$ (see Fig. 7(a),(b)). The microscopic Kirigami block is still assumed to be governed by a square lattice BHZ model. Given a different network structure the $U_{\text{eff}}$ changes. However again as a function of $\theta$ the system undergoes topological to trivial to topological phase transitions as shown in (see Fig. 7(c)-(e)). What is interesting however is that the critical points $\theta_{1c}$ and $\theta_{2c}$ are now symmetric about $\theta=\frac{\pi}{3}$. Moreover the topological transition occurs at $r(\theta)=\frac{1}{2},t(\theta)=\frac{\sqrt{3}}{2}$. Thus the transitions can be found when $\tilde{r}(\theta)=\tilde{t}(\theta)$ where $\tilde{r}(\theta)\equiv 2r(\theta)$ and $\tilde{t}(\theta)\equiv\frac{2}{\sqrt{3}}t(\theta)$. The behavior of $\tilde{r}(\theta)$ and $\tilde{t}(\theta)$ for $M=-3.5$ is shown in Fig. 7(i). The spectrum when evaluated on a ribbon geometry clearly shows the edge states in the topological regimes (see Fig. 7(f-h)). This reflects the generality of this physics to various Kirigami systems and that critical angles for transitions can themselves be tuned by nature of the Kirigami structure.
VII Discussion and Summary
Metamaterials and mechanically engineered systems herald the new era of technological devices. A host of electronic systems based on such structures have been realized. However, the interplay of quantum phases and their transitions vis-a-vis their mechanical tunability has been little explored. In this work, we show that topological phases, in particular a Chern insulator which contains chiral edge channels can be tuned to a trivial phase by such mechanical tuning. Recent experimental progress, in particular, such as Kirigami-based systems [5, 6, 10] have significantly expanded the capabilities to realize this physics.
We comment on the parameter values where our results are applicable. The quantum coherence effects of mechanical deformations etc. can be experimentally seen only when system length scales are comparable to electronic decoherence length scales. Various realistic simulations with and without molecular dynamics (MD) in graphene/hBN samples/diamond thin-films where effects of mechanical deformation on electronic structure have been studied work with Kirigami blocks where typical $L\sim 50$Å, $w\sim 20$Å and $L_{B}\sim 10$ which is at the same parameter regimes where we have performed our analysis [16, 53, 54, 55, 56]. We believe therefore realistic MD and density functional theory (DFT)-based simulations can be performed on layered topological insulating materials such as Bismuthene [57, 58, 59], HgTe [60, 61, 62] to find direct signatures of such deformation-dependent quantum transport. Experiments in graphene Kirigamis [7] and metamaterials [3, 10, 41, 63, 64] yet, are considerably larger in sizes such that one expects transport to have significant decoherence effects. These results will also be applicable to twisted bilayer graphene systems where the transport is dominated by edge currents on effective triangular networks [65, 66, 67, 68, 69]. Our results therefore motivate further experimental work in the direction of Kirigami-inspired quantum materials.
To summarize, our study places the idea of Kirigami inspired topological phases and their transitions on a concrete footing, uncovering their mechanism in a simple setting. Further studies on effects of realistic strains, gradients in deformations, strong correlations, and interactions are some of the many future directions this work opens up. Equivalent physics in alternate platforms such as photonics and phononic systems may be another interesting direction to pursue.
VIII Acknowledgement
We acknowledge fruitful discussions with G. Murthy, G. Sreejith, Diptarka Das, Arijit Kundu, Subrata Pachhal, Sudipta Dubey, Ritajit Kundu, Saikat Mondal and Soumya Sur. RS acknowledges funding from IIT Kanpur Institute Fellowship. AA acknowledges support from IITK Initiation Grant (IITK/PHY/2022010).
IX Appendix